# Fearless Stochasticity in Expectation Propagation

### v1

@@ 3.2 EP-$\bm{\eta}$
We showed in the previous section that bias is introduced into the sequence of NGD updates due to a nonlinear map from one parameterisation to the next. If the map were linear, the sequence would remain unbiased. This can be achieved by performing NGD with respect to $\eta_{i}$, the natural parameters of $\tilde{p}_{i}^{(t)}$; then, the map from one parameterisation to the next (by equating $\lambda_{i}$) is given by $\eta_{i}^{(t+1)}\circ(\eta_{i}^{(t)})^{-1}$, which is linear. The NGD direction in $L$ with respect to $\tilde{p}^{(t)}_{i}$ and $\eta_{i}$ is given by
$-\Big{[}\tilde{F}_{i}(\eta_{i})\Big{]}^{-1}\partialderivative{L}{\eta_{i}}$ $=-\alpha\partialderivative{\eta_{i}}{\mu_{i}}^{\top}\Big{(}\nabla A\big{(}\theta_{0}+\sum_{j}\lambda_{j}\big{)}-\mathbb{E}_{p_{i}(z)}[s(z)]\Big{)};$ (10)
see Appendix D for a derivation. Note that $\alpha$ plays a similar role as the NGD step size in this parameterisation, but we fix $\alpha=1$ here, and introduce an explicit NGD step size $\epsilon$. The resulting update can be expressed directly as an update in $\lambda_{i}$, by applying $(\eta_{i}^{(t)})^{-1}$ to the updated $\eta_{i}$, giving
$\lambda_{i}$ $\leftarrow\lambda_{i}-\epsilon\partialderivative{\eta_{i}}{\mu_{i}}^{\top}\Big{(}\nabla A\big{(}\theta_{0}+\sum_{j}\lambda_{j}\big{)}-\mathbb{E}_{p_{i}(z)}[s(z)]\Big{)}.$ (11)
We can use automatic differentiation to efficiently compute (11), by recognising the second term as a Jacobian-vector product (JVP) with respect to $\eta_{i}(\mu_{i})=\nabla A^{*}(\mu_{i})$. [^5] This can be computed using either forward or reverse mode automatic differentiation, as $\partialderivative{\eta_{i}}{\mu_{i}}$ is symmetric. We call the resulting procedure—which is given in Algorithm 2—EP-$\eta$, to emphasise that we have simply changed to a natural parameterisation for the NGD update of EP. The unbiased updates of EP-$\eta$ allow it to be more sample-efficient than EP by using a small number of samples per iteration. This is demonstrated using a stochastic version of the clutter problem of Minka [33] in Figure 1a.
@@ 3.3 EP-$\bm{\mu}$
(a) Expected decrease in $L$ after $100/n_{\text{samp}}$ steps. Performing e.g. 100×100$\times$100 × $1$-sample steps, or 10×10$\times$10 × $10$-sample steps, achieves a much larger decrease in $L$ than a single $100$-sample step.
(b) Magnitude of the bias in $\lambda_{i}$ after a single parallel update, averaged over all sites and dimensions. The bias of EP-$\mu$ shrinks far faster as the step size decreases than that of EP. EP-$\eta$ is always unbiased.
Figure 1: The effect of step size ($\alpha$ or $\epsilon$) and number of samples ($n_{\text{samp}}$) on different EP variants, in the clutter problem of Minka [33]. EP (naive) uses maximum likelihood estimation for the updates, and EP (debiased) uses the estimator of Xu et al. [47]. Step size corresponds to $\alpha$ for EP, and $\epsilon$ for EP-$\mu$ and EP-$\eta$. $1$-sample updates are not possible for EP, hence those traces are not visible.
The bias introduced by the updates of EP cannot be mitigated by reducing $\alpha$ (increasing the damping) without also proportionally sacrificing the amount of progress made by each update. More concretely, let the bias in dimension $d$ of $\mu_{i}^{(t+1)}$, after an update at time $t$, be defined as $E[\mu_{i}^{(t+1)}-\bar{\mu}_{i}^{(t+1)}]_{d}$, where $\bar{\mu}_{i}^{(t+1)}$ is the value of $\mu_{i}^{(t+1)}$ after a noise-free update, and expectation is taken over the sampling distributions of all parallel updates at time $t$; then, Proposition 2 below, which we prove in Appendix F, summarises the effect of decreasing $\alpha$ on EP.
@@
After update (4) is executed in parallel over $i$, as $\alpha\rightarrow 0^{+}$, both the expected decrease in $L$, and the bias $E[\mu_{i}^{(t+1)}-\bar{\mu}_{i}^{(t+1)}]_{d}$, are $O(\alpha)$ for all $d$.
Proposition 1 at the beginning of this section states that update (4) can be viewed as performing an NGD step with respect to $\mu_{i}$ with a step size of $\alpha^{-1}$. However, changing $\alpha$ also has the effect of changing the definition of $\tilde{p}_{i}^{(t)}$. It is then natural to wonder what happens if we fix $\alpha=1$ and introduce an explicit step size $\epsilon$ for NGD, resulting in the update
$\mu_{i}$ $\leftarrow\mu_{i}-\epsilon\Big{(}\nabla A\big{(}\theta_{0}+\sum_{j}\lambda_{j}\big{)}-\mathbb{E}_{p_{i}(z)}[s(z)]\Big{)};$ (12)
see Appendix E for a derivation. It turns out that in doing so, when we decrease $\epsilon$, the bias shrinks far faster than the expected decrease in $L$.
@@
After update (12) is executed in parallel over $i$, as $\epsilon\rightarrow 0^{+}$, the expected decrease in $L$ is $O(\epsilon)$, and the bias $E[\mu_{i}^{(t+1)}-\bar{\mu}_{i}^{(t+1)}]_{d}$ is $O(\epsilon^{2})$, for all $d$.
Proposition 3, which we prove in Appendix F, tells us that by using update (12), we can reduce the bias to arbitrarily small levels while still making progress in decreasing $L$. Update (12) can also be expressed directly as an update in $\lambda_{i}$ by applying $(\eta^{(t)})^{-1}\circ\nabla A^{*}$, giving
$\lambda_{i}$ $\leftarrow\nabla A^{*}\Big{(}(1-\epsilon)\nabla A\big{(}\theta_{0}+\sum_{j}\lambda_{j}\big{)}+\epsilon\mathbb{E}_{p_{i}(z)}[s(z)]\Big{)}-\theta_{0}-\sum_{j\neq i}\lambda_{j},$ (13)
which has the simple interpretation of performing EP, but with damping of the moments instead of the site parameters. We call this variant EP-$\mu$, again to indicate that we are simply performing the NGD update of EP in the mean parameter space. [^6] EP too is performing NGD in mean parameters, but using a step size that also affects the distribution $\tilde{p}_{i}$. The resulting procedure is also given by Algorithm 2.
@@
$\mathcal{F}$, $\theta_{0}$, $\{\beta_{i}\}_{i}$, $\{\ell_{i}(z)\}_{i}$, $\{\lambda_{i}\}_{i}$, $n_{\text{inner}}$, $\epsilon$
@@
     $\theta_{i}\leftarrow\theta_{0}+\sum_{j}\lambda_{j}$ for $i=1$ to $m$
@@
The computational cost of EP-$\mu$ is lower than that of EP-$\eta$ because it does not require JVPs through $\nabla A^{*}(.)$. [^7] See Appendix G for a detailed discussion of the computational costs of EP-$\eta$ and EP-$\mu$. The drawback is that the updates of EP-$\mu$ do still retain some bias; however, we find that the bias of EP-$\mu$ typically has negligible impact on its performance relative to EP-$\eta$. This is evident in the clutter problem of Minka [33], as shown in Figure 1a, as well as in the larger scale experiments of Section 4. Figure 1b illustrates how quickly the bias in $\lambda_{i}$ shrinks as the step size of EP-$\mu$ is decreased; note that the results of this subsection are stated in terms of $\mu_{i}$, but it is straightforward to show that equivalent results also hold for $\lambda_{i}$ using Taylor series arguments.
@@ 3.4 Related work
EP was combined with MC moment estimates by Xu et al. [47] in a method called sampling via moment sharing (SMS), and later by Vehtari et al. [44] in the context of hierarchical models. In both works, when $\mathcal{F}$ was multivariate normal—arguably the most useful case—the authors used an estimator for the updates that is unbiased when $\tilde{p}_{i}$ is also multivariate normal. Although this estimator is only approximately unbiased in general, it can help to stabilise the updates significantly. Even so, it is necessary to use a relatively large number of samples for the updates, with several hundred or more being typical. Using many samples per update is inefficient, as the update direction can change from one iteration to the next. Furthermore, such estimators typically rely on a known number of independent samples; samples drawn using Markov chain Monte Carlo (MCMC) methods are generally autocorrelated, necessitating the use of sample thinning before the estimators can be applied. This adds another hyperparameter to the procedure—the thinning ratio—and it is also inefficient due to the discarding of samples. The practitioner is required to choose the number of samples, the thinning ratio, and the amount of damping, all of which affect the accuracy, stability, and computational efficiency of the procedure. There is no easy way to choose these a priori, forcing the practitioner to either set them conservatively (favouring accuracy and stability), or to find appropriate settings by trial and error, both of which are likely to expend time and computation unnecessarily.
@@
The stochastic natural gradient EP (SNEP) method of Hasenclever et al. [14] optimises the same variational objective as EP, but performs stochastic natural gradient descent with respect to a mean parameterisation (under $\mathcal{F}$) of the site potentials in order to perform the inner minimisation. Its natural gradient updates are unbiased in the presence of MC noise, allowing them to be performed with as little as one sample, without relying on $\mathcal{F}$-specific debiasing estimators, or the sample thinning that would entail. The authors found that in some settings, accurate point estimates can be obtained fairly rapidly with SNEP. However, we find that SNEP typically converges far slower than EP, consistent with findings in Vehtari et al. [44]. In SNEP the site potentials are treated as bona fide distributions for the purposes of optimisation, and can be seen as playing the role of surrogate distributions for NGD [42]. The results of this section showed that EP, EP-$\eta$ and EP-$\mu$ are closely related to SNEP; they too can be viewed as performing NGD of $L$ with respect to surrogate distributions ($\tilde{p}_{i}$), but with distributions that are more closely matched with those being optimised. This allows them to retain the same benefits as SNEP, but with faster convergence.
@@ 4 Evaluation
In this section we demonstrate the efficacy of EP-$\eta$ and EP-$\mu$ on a variety of probabilistic inference tasks. In each experiment, the task was to perform approximate Bayesian inference of unobserved parameters $z$, given some observed data $\mathcal{D}$. All of the models in these experiments followed the same general structure, consisting of a minimal exponential family prior over $z$, and a partition of the data $\{\mathcal{D}_{i}\}_{i=1}^{m}$, where each block $\mathcal{D}_{i}$ depends on both $z$ and an additional vector of local latent variables $w_{i}$ that also depend on $z$; that is, the joint density has structure
$p_{0}(z)\prod_{i=1}^{m}p(w_{i}\mid z)p(\mathcal{D}_{i}\mid w_{i}),$ (14)
where $p_{0}\in\mathcal{F}$. This is shown graphically in Appendix H. To perform approximate inference of $z$, we first define $\ell_{i}(z)$ as the log likelihood of $\mathcal{D}_{i}$ given $z$, with $w_{i}$ marginalised out; that is,
@@
Then, given $p_{0}(z)$ and $\{\ell_{i}(z)\}_{i}$, we define $p^{*}(z)$ as (1), and proceed to find an approximation $p(z)\approx p^{*}(z)$ using the methods described in earlier sections. Note that sampling from the tilted distribution $p_{i}(z)$ requires jointly sampling over $z$ and $w_{i}$ and taking the marginal. EP is a particularly appealing framework for performing inference in this setting, as the dimensionality of the sampled distributions is constant with respect to $m$, mitigating the curse of dimensionality experienced by conventional MCMC approaches [44]. In our experiments we used NUTS [24] to perform the sampling, consistent with prior work [44, 47]. In each experiment we compared EP-$\eta$ and EP-$\mu$ with EP, in the manner used by Xu et al. [47], Vehtari et al. [44]. Where $\mathcal{F}$ was MVN, we also compared to SNEP [14]. For the models with NIW $\mathcal{F}$, we were unable to find an initialisation for SNEP that would allow us to perform a meaningful comparison; we discuss this more in Appendix H.
@@
To evaluate the performance of the different variants, we monitored KL divergence to an estimate of the optimum, obtained by running EP with a large number of samples. We used 500 different hyperparameter settings for each variant, chosen using random search, and repeated each run 5 times using different random seeds for NUTS. All runs of EP-$\eta$ and EP-$\mu$ were performed with $n_{\text{samp}}=1$ and $n_{\text{inner}}=1$. The results of these experiments are summarised in Figure 2
(a) Hierarchical logistic regression: MVN prior
(b) Hierarchical logistic regression: NIW prior
(c) Cosmic radiation model
(d) Neural response model
@@
Figure 2: Pareto frontier showing the number of NUTS steps ($x$-axis) against the KL divergence from $p$ to an estimate of the optimum ($y$-axis). Each point on the plot marks the lowest average KL divergence attained by any hyperparameter setting by that step count. Error bars mark the full range of values for the marked hyperparameter setting across 5 random seeds.
@@
We see that EP-$\eta$ and EP-$\mu$ can typically reach a given level of accuracy (distance from optimum) faster than EP, often significantly so. We stress that the frontiers for EP-$\eta$ and EP-$\mu$ are traced out with $n_{\text{samp}}$ and $n_{\text{inner}}$ each fixed to $1$, with no sample thinning, varying only the step size hyperparameter $\epsilon$, with higher accuracy/cost regions of the frontier corresponding to lower values of $\epsilon$. In contrast, to trace out the frontier of EP we must jointly vary $\alpha$, $n_{\text{samp}}$, and the thinning ratio; see Appendix I for examples of this effect. The performances of EP-$\eta$ and EP-$\mu$ are very similar. We found SNEP to be significantly slower than the other methods, consistent with findings by Vehtari et al. [44]. Pareto frontiers with respect to wall-clock time can be found in Appendix J. We now provide a brief overview of individual experiments, with further details given in Appendix H.
4.1 Hierarchical logistic regression
Hierarchical logistic regression is used to perform binary classification in a number of groups when the extent to which data should be pooled across groups is unknown [12]. Each of the $m$ groups has its own covariates and binary response variables, collectively denoted by $\mathcal{D}_{i}$, as well as regression coefficients $w_{i}$. The groups are related through prior parameters $z$, which parameterise the group-level coefficient prior densities $p(w_{i}\mid z)$. We performed two experiments in this setting.
Synthetic data with MVN prior
In this experiment, the group-level prior density $p(w_{i}\mid z)$ was a product of independent normal densities, with $p_{0}\in\mathcal{F}$ being a multivariate normal (MVN) prior over the means and log-variances. We used the same synthetic data generation as Vehtari et al. [44], which was designed to be challenging for EP. In this experiment $z\in\mathbb{R}^{8}$, $w_{i}\in\mathbb{R}^{4}\ \forall\ i$, $m=16$, and each of the $m$ regression tasks had $n=20$ observations. Results are shown in Figure 2a.
Political survey data with NIW prior
@@ 4.1 Hierarchical logistic regression
In this experiment, each of $m=50$ nested regression tasks corresponded to political survey responses from a particular US state, with $7$ predictor variables corresponding to characteristics of a given survey respondent, so that $w_{i}\in\mathbb{R}^{7}$ for $i=1,\ldots,m$. The state-level prior density $p(w_{i}\mid z)$ was that of a MVN distribution, with $p_{0}\in\mathcal{F}$ being a normal-inverse-Wishart (NIW) prior on the mean and covariance matrix. This structure allows for correlation of regression coefficients across states. The data, taken from the 2018 Cooperative Congressional Election Study, related to support for allowing employers to decline coverage of abortions in insurance plans [29, 39]. We truncated the data slightly so that each state contained $n=97$ survey responses. Results are shown in Figure 2b.
4.2 Cosmic radiation model

### v2

@@ 3.2 EP-$\bm{\eta}$
We showed in the previous section that bias is introduced into the sequence of NGD updates due to a nonlinear map from one parameterisation to the next. If the map were affine, the sequence would remain unbiased. This can be achieved by performing NGD with respect to $\eta_{i}$, the natural parameters of $\smash{\cramped{\tilde{p}_{i}}^{(t)}}$. Then, the map from one parameterisation to the next (by equating $\lambda_{i}$) is given by $\smash{\cramped{\eta_{i}}^{(t+1)}\circ\cramped{(\cramped{\eta_{i}}^{(t)})}^{-1}}$, which is linear. The NGD direction in $L$ with respect to $\smash{\cramped{\tilde{p}_{i}}^{(t)}}$ and $\eta_{i}$ is given by
$-\Big{[}\tilde{F}_{i}(\eta_{i})\Big{]}^{-1}\partialderivative{L}{\eta_{i}}$ $=-\alpha\partialderivative{\eta_{i}}{\mu_{i}}\Big{(}\nabla A\big{(}\eta_{0}+\sum_{j}\lambda_{j}\big{)}-\mathbb{E}_{p_{i}(z)}[s(z)]\Big{)}.$ (10)
See Appendix E for a derivation. Note that $\alpha$ plays a similar role as the NGD step size in this parameterisation, but it also affects $\smash{\cramped{\tilde{p}_{i}}^{(t)}}$, and so to obtain a more faithful NGD interpretation we fix $\alpha=1$ and introduce an explicit NGD step size $\epsilon$ to decouple from the effect on $\smash{\cramped{\tilde{p}_{i}}^{(t)}}$. The resulting update can be expressed directly as an update in $\lambda_{i}$, by applying $\smash{\cramped{(\cramped{\eta_{i}}^{(t)})}^{-1}}$ to the updated $\eta_{i}$, giving
$\lambda_{i}$ $\leftarrow\lambda_{i}-\epsilon\partialderivative{\eta_{i}}{\mu_{i}}\Big{(}\nabla A\big{(}\eta_{0}+\sum_{j}\lambda_{j}\big{)}-\mathbb{E}_{p_{i}(z)}[s(z)]\Big{)}.$ (11)
We can use automatic differentiation to efficiently compute (11), by recognising the second term as a Jacobian-vector product (JVP) with respect to $\eta_{i}(\mu_{i})=\nabla A^{*}(\mu_{i})$. [^5] This can be computed using either forward or reverse mode automatic differentiation, as $\partialderivative{\eta_{i}}{\mu_{i}}$ is symmetric. In summary, by performing NGD with respect to natural parameters of $\smash{\cramped{\tilde{p}_{i}}^{(t)}}$, instead of mean parameters, the resulting sequence of updates is unbiased in $\lambda_{i}$. We call the resulting procedure—which is given in Algorithm 2—EP-$\eta$, to emphasise that we have simply changed the NGD parameterisation of EP. The unbiased updates of EP-$\eta$ allow it to be more sample-efficient than EP by using a smaller number of samples per iteration. The left panel of Figure 1 demonstrates this effect in a stochastic version of the clutter problem [37].
@@ 3.3 EP-$\bm{\mu}$
Figure 1: The effect of step size ($\alpha$ or $\epsilon$) and number of MC samples ($n_{\text{samp}}$) on different EP variants in a stochastic version of the clutter problem of Minka [37]. EP (naive) uses maximum likelihood estimation for the updates, and EP (debiased) uses the estimator of Xu et al. [52]. Step size corresponds to $\alpha$ for EP, and $\epsilon$ for EP-$\mu$ and EP-$\eta$. Only EP-$\eta$ and EP-$\mu$ can perform $1$-sample updates, hence the other traces are not visible. The left panel shows the expected decrease in $L$ after $100/n_{\text{samp}}$ steps. Performing e.g. 100×100$\times$100 × $1$-sample steps, or 10×10$\times$10 × $10$-sample steps, achieves a much larger decrease in $L$ than a single $100$-sample step. The right panel shows the magnitude of the bias in $\lambda_{i}$ after a single parallel update, averaged over all sites and dimensions. The bias of EP-$\mu$ shrinks far faster as the step size decreases than that of EP. EP-$\eta$ is always unbiased and so is not visible.
The bias introduced by the updates of EP cannot be mitigated by reducing $\alpha$ (increasing the damping) without also proportionally sacrificing the amount of progress made by each update. More concretely, let the bias in dimension $d$ of $\smash{\cramped{\mu_{i}}^{(t+1)}}$, after an update at time $t$, be defined as $\smash{\mathbb{E}[\cramped{\mu_{i}}^{(t+1)}-\cramped{\bar{\mu}_{i}}^{(t+1)}]_{d}}$, where $\smash{\cramped{\bar{\mu}_{i}}^{(t+1)}}$ is the value of $\smash{\cramped{\mu_{i}}^{(t+1)}}$ after a noise-free update, and expectation is taken over the sampling distributions of all parallel updates at time $t$. Then, Proposition 2 below, which we prove in Appendix G, summarises the effect of decreasing $\alpha$ on EP.
@@
After update (4) is executed in parallel over $i$, as $\alpha\rightarrow 0^{+}$, both the expected decrease in $L$, and the bias $\smash{\mathbb{E}[\cramped{\mu_{i}}^{(t+1)}-\cramped{\bar{\mu}_{i}}^{(t+1)}]_{d}}$, are $O(\alpha)$ for all $d$.
Proposition 1 at the beginning of this section states that update (4) can be viewed as performing an NGD step with respect to $\mu_{i}$ with a step size of $\alpha^{-1}$. However, changing $\alpha$ also has the effect of changing the definition of $\smash{\cramped{\tilde{p}_{i}}^{(t)}}$. It is then natural to wonder what happens if we fix $\alpha=1$ and introduce an explicit step size for NGD, $\epsilon$, resulting in the update
$\mu_{i}$ $\leftarrow\mu_{i}-\epsilon\Big{(}\nabla A\big{(}\eta_{0}+\sum_{j}\lambda_{j}\big{)}-\mathbb{E}_{p_{i}(z)}[s(z)]\Big{)}.$ (12)
See Appendix F for a derivation. It turns out that in doing so, when we decrease $\epsilon$, the bias shrinks far faster than the expected decrease in $L$.
@@
After update (12) is executed in parallel over $i$, as $\epsilon\rightarrow 0^{+}$, the expected decrease in $L$ is $O(\epsilon)$, and the bias $\smash{\mathbb{E}\left[\cramped{\mu_{i}}^{(t+1)}-\cramped{\bar{\mu}_{i}}^{(t+1)}\right]_{d}}$ is $O(\epsilon^{2})$, for all $d$.
Proposition 3, which we prove in Appendix G, tells us that by using update (12), we can reduce the bias to arbitrarily small levels while still making progress in decreasing $L$. Update (12) can also be expressed directly as an update in $\lambda_{i}$ by applying $\smash{\cramped{(\eta^{(t)})}^{-1}\circ\nabla A^{*}}$, giving
$\lambda_{i}$ $\leftarrow\nabla A^{*}\Big{(}(1-\epsilon)\nabla A\big{(}\eta_{0}+\sum_{j}\lambda_{j}\big{)}+\epsilon\mathbb{E}_{p_{i}(z)}[s(z)]\Big{)}-\eta_{0}-\sum_{j\neq i}\lambda_{j},$ (13)
which has the simple interpretation of performing EP, but with damping of the moments instead of the site (natural) parameters. In summary, if we perform NGD with respect to the same mean parameterisation as EP, but treat the NGD step size as a free parameter, $\epsilon$, we can obtain updates that are approximately unbiased while still making progress. We call this variant EP-$\mu$, again to indicate that we are simply performing the NGD update of EP in the mean parameter space. [^6] EP too is performing NGD in mean parameters, but using a “step size” that also affects the distribution $\tilde{p}_{i}$. The resulting procedure is also given by Algorithm 2.
@@
$\mathcal{F}$, $\eta_{0}$, $\{\beta_{i}\}_{i}$, $\{\ell_{i}(z)\}_{i}$, $\{\lambda_{i}\}_{i}$, $n_{\text{inner}}$, $\epsilon$
@@
     $\smash{\theta\leftarrow\eta_{0}+\cramped{\sum}_{j}\lambda_{j}}$
@@
The computational cost of EP-$\mu$ is lower than that of EP-$\eta$ because it does not require JVPs through $\nabla A^{*}(.)$. [^7] See Appendix I for a detailed discussion of the computational costs of EP-$\eta$ and EP-$\mu$. The drawback is that the updates of EP-$\mu$ do still retain some bias, however, we find that the bias of EP-$\mu$ typically has negligible impact on its performance relative to EP-$\eta$. This is evident in the clutter problem of Minka [37], as demonstrated in the left panel of Figure 1, as well as in the larger scale experiments of Section 4. The right panel of Figure 1 illustrates how quickly the bias in $\lambda_{i}$ shrinks as the step size of EP-$\mu$ is decreased. Note that the results of this subsection are stated in terms of bias in $\mu_{i}$, but it is straightforward to show that equivalent results also hold for $\lambda_{i}$ using Taylor series arguments.
@@ 3.4 Related work
The link between natural gradients and moment matching in an exponential family is well known [6], but the connection with EP shown here – which relies on identification of the distribution $\smash{\cramped{\tilde{p}_{i}}^{(t)}}$ and parameterisation $\smash{\cramped{\eta_{i}}^{(t)}}$ – is new, to the best of our knowledge. Bui et al. [11] showed that the updates of (power) EP are equivalent to performing NGD in a local variational free energy objective when taking the limit of the power parameter as $\beta_{i}\rightarrow\infty$, and Wilkinson et al. [51] showed that this extends to the (global) variational free energy objective for models with a certain structure. Our result is different in that it relates to NGD of the EP variational objective, and applies for all values of $\beta_{i}$.
@@
EP was combined with Markov chain Monte Carlo (MCMC) moment estimates by Xu et al. [52] in a method called sampling via moment sharing (SMS), and later by Vehtari et al. [48] in the context of hierarchical models. In both works, when $\mathcal{F}$ was multivariate normal (MVN) —arguably the most useful case—the authors used an estimator for the updates that is unbiased when $\tilde{p}_{i}$ is also MVN. Although this estimator is only approximately unbiased in general, it can help to stabilise the updates significantly. Even so, it is necessary to use a relatively large number of samples for the updates, with several hundred or more being typical. Using many samples per update is inefficient, as the update direction can change from one iteration to the next. Furthermore, such estimators typically rely on a known number of independent samples. Samples drawn using MCMC methods are generally autocorrelated, necessitating the use of sample thinning before the estimators can be applied. This adds another hyperparameter to the procedure—the thinning ratio—and it is also inefficient due to the discarding of samples. The practitioner is required to choose the number of samples, the thinning ratio, and the amount of damping, all of which affect the accuracy, stability, and computational efficiency of the procedure. There is no easy way to choose these a priori, forcing the practitioner to either set them conservatively (favouring accuracy and stability), or to find appropriate settings by trial and error, both of which are likely to expend time and computation unnecessarily.
@@
The stochastic natural gradient EP (SNEP) method of Hasenclever et al. [17], which is closely related to our work, also optimises the EP variational objective using a form of NGD. The SNEP updates are unbiased in the presence of MC noise, allowing them to be performed with as little as one sample, without relying on $\mathcal{F}$-specific debiasing estimators, or the sample thinning that would entail. In SNEP, NGD is performed with respect to mean parameters of the site potentials, which are treated as bona fide distributions in $\mathcal{F}$. In contrast, we showed that EP can already be viewed as performing NGD, but with respect to the distributions $\smash{\{\cramped{\tilde{p}_{i}}^{(t)}\}_{i}}$, and our new variants, EP-$\eta$ and EP-$\mu$, are able to gain the same advantages as SNEP, but using the same distributions for NGD as EP. Hasenclever et al. [17] showed that in some settings, SNEP can obtain accurate point estimates fairly rapidly. However, we find that it typically converges far slower than both EP and our new variants, consistent with findings in Vehtari et al. [48]. We argue that this is because the site potentials (when considered as distributions) bear little resemblance with the distributions that are ultimately being optimised, and so their geometry is largely irrelevant for the optimisation of $L$. In contrast, the geometry of $\smash{\cramped{p_{i}}^{(t)}}$ is closely related to that of $L$, as we show in Appendix H.
@@ 4 Evaluation
In this section we demonstrate the efficacy of EP-$\eta$ and EP-$\mu$ on a variety of probabilistic inference tasks. In each experiment, the task was to perform approximate Bayesian inference of unobserved parameters $z$, given some observed data $\mathcal{D}$. All of the models in these experiments followed the same general structure, consisting of a minimal exponential family prior over $z$, $p_{0}$, and a partition of the data $\{\mathcal{D}_{i}\}_{i}^{m}$, where each block $\mathcal{D}_{i}$ depends on both $z$ and an additional vector of local latent variables $w_{i}$ that also depend on $z$. That is, the joint density has the form
$p_{0}(z)\prod_{i}p(w_{i}\mid z)p(\mathcal{D}_{i}\mid w_{i},z),$ (14)
where $p_{0}\in\mathcal{F}$. This structure is shown graphically in Appendix J. To perform approximate inference of $z$, we first define $\ell_{i}(z)$ as the log likelihood of $\mathcal{D}_{i}$ given $z$, with $w_{i}$ marginalised out. That is,
@@
Then, given $p_{0}(z)$ and $\{\ell_{i}(z)\}_{i}$, we define $p^{*}(z)$ as (1), and proceed to find an approximation $p(z)\approx p^{*}(z)$ using the methods described in earlier sections. Note that sampling from the tilted distribution $p_{i}(z)$ requires jointly sampling over $z$ and $w_{i}$ and taking the marginal. EP is a particularly appealing framework for performing inference in this setting, as the dimensionality of the sampled distributions is constant with respect to $m$, mitigating the curse of dimensionality experienced by conventional MCMC approaches [48]. In our experiments we used NUTS [27] to perform the sampling, consistent with prior work [48, 52]. In each experiment we compared EP-$\eta$ and EP-$\mu$ with EP, in the manner used by Xu et al. [52] and Vehtari et al. [48]. When $\mathcal{F}$ was MVN we also compared to SNEP [17]. For the models with normal-inverse Wishart (NIW) $\mathcal{F}$, we were unable to find an initialisation for SNEP that would allow us to perform a meaningful comparison. We discuss this point further in Appendix J.
@@
To evaluate the performance of the different variants, we monitored KL divergence to an estimate of the optimum, obtained by running EP with a large number of samples. We used 500 different hyperparameter settings for each variant, chosen using random search, and repeated each run 5 times using different random seeds for NUTS. All runs of EP-$\eta$ and EP-$\mu$ were performed with $n_{\text{samp}}=1$ and $n_{\text{inner}}=1$. The results of these experiments are summarised by the Pareto frontiers in Figure 2
@@
Figure 2: Pareto frontiers showing the number of NUTS steps ($x$-axis) against the KL divergence from $p$ to an estimate of the optimum ($y$-axis). Each point on the plot marks the lowest average KL divergence attained by any hyperparameter setting by that step count. Error bars mark the full range of values for the marked hyperparameter setting across 5 random seeds.
@@
We see that EP-$\eta$ and EP-$\mu$ can typically reach a given level of accuracy (distance from an estimate of the optimum) faster than EP, often significantly so. We stress that the frontiers for EP-$\eta$ and EP-$\mu$ are traced out with $n_{\text{samp}}$ and $n_{\text{inner}}$ each fixed to $1$, with no sample thinning, varying only the step size hyperparameter $\epsilon$, with higher accuracy/cost regions of the frontier corresponding to lower values of $\epsilon$. In contrast, to trace out the frontier of EP we must jointly vary $\alpha$, $n_{\text{samp}}$, and the thinning ratio. See Appendix K for examples of this effect. The performances of EP-$\eta$ and EP-$\mu$ are very similar. We found SNEP to be significantly slower than the other methods, consistent with findings by Vehtari et al. [48]. Pareto frontiers with respect to wall-clock time can be found in Appendix L. Code for these experiments can be found at https://github.com/cambridge-mlg/fearless-ep. We now provide a brief overview of individual experiments, with further details given in Appendix J.
Hierarchical logistic regression with MVN prior
Hierarchical logistic regression (HLR) is used to perform binary classification in a number of groups when the extent to which data should be pooled across groups is unknown [15]. We performed approximate Bayesian inference in a HLR model using a MVN prior over the group-level parameters. We applied this model to synthetic data, generated using the same procedure as Vehtari et al. [48], which was designed to be challenging for EP. In this experiment $z\in\mathbb{R}^{8}$, $w_{i}\in\mathbb{R}^{4}\ \forall\ i$, and each of the $m=16$ groups had $n=20$ observations. In the upper-left panel of Figure 2 we see that EP-$\eta$ and EP-$\mu$ typically reach a given level of accuracy faster than EP and SNEP.
Hierarchical logistic regression with NIW prior
@@
We also performed approximate Bayesian inference in a similar HLR model, but using a normal-inverse Wishart (NIW) prior over group-level parameters, allowing for correlation of regression coefficients within groups. We applied this model to political survey data, where each of $m=50$ groups corresponded to responses from a particular US state, with $7$ predictor variables corresponding to characteristics of a given survey respondent, so that $w_{i}\in\mathbb{R}^{7}\;\forall\;i$. There were $n=97$ survey respondents per state. The data, taken from the 2018 Cooperative Congressional Election Study, related to support for allowing employers to decline coverage of abortions in insurance plans [33, 43]. The upper-right panel of Figure 2 shows that EP-$\eta$ and EP-$\mu$ consistently reach a given level of accuracy significantly faster than EP.
Cosmic radiation model